### v1

@@ E.1 Dominant Label Ratio
E.2 Dirichlet partitioning
Dirichlet partitioning is probably the most commonly used method to federate data in distributed computing research solans2024non. It leverages the Dirichlet Distribution minka2000dirichlet to allocate samples among clients, effectively simulating non-IID scenarios. The concentration parameter, $\alpha$, determines the level of data heterogeneity: lower values result in more imbalanced distributions, while higher values produce distributions closer to IID conditions. This approach enables the creation of diverse non-IID datasets, facilitating the evaluation of FL algorithms under varying levels of data imbalance.
@@ E.3 Custom sharding method
The sharding technique zhang2021federated generates non-IID datasets by first sorting data based on class labels and dividing it into smaller, equally sized subsets known as shards. These shards are then distributed to clients, with each client receiving a fixed number. As each shard contains data from only a few classes, clients end up with datasets representing only a limited portion of the overall class distribution. This approach creates a highly skewed, non-IID distribution, where clients predominantly have samples from a small number of classes, without any overlap in class representation across clients from different shards.
@@ F.2 Cyclic-and-reverse order
Appendix G Extended Systematic Analysis
@@
G.1 Studying performance for other data patterns
Table 5: PG and global accuracy (reported median of the last five rounds across all runs) of SFL for other data patterns, supplementary to Table 1. The results presented here are for cyclic order with $\phi=10$. For the IID, it presents in the parentheses the PG and the global accuracy, respectively.
Performance Gap $\downarrow$ Global Accuracy $\uparrow$
@@ G.1 Studying performance for other data patterns
CIFAR-10
MobileNet (10, 80) $\mathbf{36}$ $\mathbf{\pm 1}$ $43.7$ $\pm 0.5$ $45.5$ $\pm 1$ $\mathbf{50}$ $\mathbf{\pm 0.4}$ $\mathbf{50}$ $\mathbf{\pm 0.5}$ $40$ $\pm 0.6$
ReNet101 (15, 68) $\mathbf{35}$ $\mathbf{\pm 3}$ $53$ $\pm 2$ $55$ $\pm 3$ $\mathbf{42}$ $\mathbf{\pm 1.2}$ $33$ $\pm 0.6$ $27$ $\pm 1$
TinyImageNet $(\mathbf{L=200})$
MobileNet (40, 50) $62$ $\pm 1.6$ $\mathbf{46}$ $\mathbf{\pm 1}$ $63$ $\pm 0.7$ $27$ $\pm 0.3$ $\mathbf{45}$ $\mathbf{\pm 0.3}$ $31$ $\pm 1$
@@ G.4 Studying the per-position Accuracy
Per-position Accuracy. The position refers to the order in which client $c_{l}$ with dominant label $l$ is processed at the server. In practice, position-$k$ accuracy, where $k\in\{1\ldotp\ldotp L\}$, is the accuracy of the labels that are dominant in clients that are processed at position $k$ (in the processing cycle of length $L$). To ensure that the results are independent of the type of label, the order of the labels is randomly allocated in every experiment, and then the median value from all repeated experiments is computed to give the per-position accuracy. This definition is tailored to the cyclic order and DL partitioning, but can be extended to other types, too.
In the following, we present supplemental results to the ones presented in Section 3. Note that we incorporate three different models, and for each of them, we use three different cut-layer levels (i.e., shallow-cut, mid-cut, deep-cut). In detail, the definition of these cuts, respectively, for each level, is: (i) MobileNet has $26$ layers in total and uses cuts $4,15$, and $23$, (ii) ResNet101 has $37$ layers in total and uses cuts $2,10$, and $33$, (iii) ResNet18 has $12$ layers in total and uses cuts $2,6$, and $10$. Note that for the ResNet models we consider that each residual block is a single layer, like in tirana2024workflow. Moreover, recall that the cut-layer selection follows the bottleneck policy as discussed in Appendix D.2.
@@
G.5 Scatter plot
@@
Figure 16 presents an extended version of the scatter plot shown in Figure 3, which depicts the global accuracy and the PG for different cut layers (in SFL), and for SplitFedv1 thapa2022splitfed, a version of SL. It has been shown that its performance is the same as that of FL thapa2022splitfed, and hence, we denote it by FL. Here we utilize the same cut-layers as in Appendix G.4. We see that in terms of accuracy, SFL performs better than FL under IID data. Still, as data heterogeneity increases, FL performs better. While these results concern cyclic order and DL partition, results on random order and Dirichlet partition are provided in hanconvergence. Similar are the results for CIFAR-100, even though the performance is not affected as much when the data heterogeneity increases. This is a side-effect of our earlier observation made in Appendix G.4 (i.e., lower $L$, present higher intraCF).
@@ Appendix H Extended Numerical Evaluation
Performance Gap $\downarrow$ Global Accuracy $\uparrow$
@@
SFL+Hydra $\mathbf{8.3_{\pm 0.6}}$ $\mathbf{8.3_{\pm 0.5}}$ $\mathbf{10_{\pm 0.4}}$ $\mathbf{9.3_{\pm 0.6}}$ $\mathbf{81.3_{\pm 0.1}}$ $\mathbf{81_{\pm 0.1}}$ $\mathbf{81_{\pm 0.1}}$ $\mathbf{82_{\pm 0.3}}$
SFL $24_{\pm 0.3}$ $24_{\pm 0.4}$ $21.8_{\pm 0.2}$ $23.8_{\pm 0.7}$ $7.4_{\pm 0.4} 77$$7.7_{\pm 0.4}$ $8.2_{\pm 0.5} 77$$7.7_{\pm 0.6}$
@@
SFL+Hydra $\mathbf{13_{\pm 0.9}}$ $\mathbf{13_{\pm 1.5}}$ $\mathbf{15_{\pm 0.6}}$ $\mathbf{12_{\pm 0.8}}$ $\mathbf{75.2_{\pm 1.4}}$ $\mathbf{74.6_{\pm 0.8}}$ $\mathbf{73.4_{\pm 1.5}}$ $\mathbf{66_{\pm 0.9}}$
SFL $50_{\pm 3}$ $46_{\pm 3}$ $42_{\pm 1}$ $54.9_{\pm 1}$ $56_{\pm 5}$ $7_{\pm 3} 66$$6.2_{\pm 1.6}$ $59.7_{\pm 2}$
@@
Performance Gap $\downarrow$ Global Accuracy $\uparrow$
@@
SFL + Hydra $\mathbf{22_{\pm 2}}$ $\mathbf{25_{\pm 2}}$ $\mathbf{30_{\pm 2}}$ $\mathbf{45_{\pm 1}}$ $\mathbf{45.7_{\pm 1}}$ $\mathbf{39_{\pm 0.4}}$
SFL $35_{\pm 3}$ $3_{\pm 2} 55$$5_{\pm 3}$ $42_{\pm 1.2}$ $33_{\pm 0.6}$ $27_{\pm 1}$
@@ H.1 SFL and Hydra
H.2 Other types of ordering
@@ H.3 Results with other data partitioning methods
Performance Gap $\downarrow$ Global Accuracy $\uparrow$
@@
SFL+Hydra $\mathbf{13.4_{\pm 1}}$ $\mathbf{12_{\pm 1}}$ $\mathbf{26.3_{\pm 0.66}}$ $\mathbf{66.5}$$\mathbf{\pm 0.1}$ $\mathbf{74.6{\pm 0.1}}$ $\mathbf{45_{\pm 0.1}}$
SplitFedV1/FL $15.7_{\pm 1.5}$ $15{\pm 0.4}$ $34.46_{\pm 0.3}$ $55_{\pm 0.2}$ $63.4_{\pm 0.1}$ $32_{\pm 0.19}$
SplitFedV3 $2.6_{\pm 0.8} 22$$2_{\pm 1}$ $47.3_{\pm 0.6}$ $43_{\pm 0.3}$ $51.4_{\pm 0.1}$ $25.9_{\pm 0.1}$
MultiHead $36.6_{\pm 0.9}$ $30.5_{\pm 1}$ $59.9_{\pm 0.3}$ $47_{\pm 0.3}$ $55.7_{\pm 0.46}$ $25.4_{\pm 0.18}$
SplitNN $44_{\pm 1}$ $24_{\pm 0.7}$ $37_{\pm 0.8}$ $53_{\pm 1}$ $72_{\pm 0.6}$ $42_{\pm 0.2}$
@@
SFL + Hydra $\mathbf{16.6_{\pm 3.4}}$ $\mathbf{17_{\pm 2.4}}$ $\mathbf{33_{\pm 1}}$ $\mathbf{49.7_{\pm 1.4}}$ $\mathbf{51.9_{\pm 0.9}}$ $\mathbf{33_{\pm 0.9}}$
SplitFedV1/FL $22.8_{\pm 2.4}$ $24.1_{\pm 4}$ $43_{\pm 1.6}$ $29.7_{\pm 1.8}$ $39_{\pm 1.1}$ $23_{\pm 1.6}$
SplitFedV3 $40_{\pm 3}$ $27.1_{\pm 1}$ $19.1_{\pm 1.6}$ $35_{\pm 2}$
MultiHead $33_{\pm 2}$ $28_{\pm 4}$ $29_{\pm 1}$ $40_{\pm 0.9}$
SplitNN $73.5_{\pm 0.8}$ $44_{\pm 2}$ $79_{\pm 1.6}$ $23.4_{\pm 1}$ $48.7_{\pm 1.8}$ $16.7_{\pm 0.4}$
H.4 Hydra compared to baseline approaches
@@ H.5 Weight Regularization
H.6 Ablation study on the number of heads.
In this section, we first provide details about the definition of the superclasses in CIFAR-10 for the results presented in Table 4. In particular, similar in spirit of bai2021clustering, we group labels according to their semantic relationships. In detail, for $G=5$, we have defined the groups/superclasses: land vehicles: automobile, truck, things found in the sky: airplane, bird, water vehicles: ship, domestic animals: cat, dog, wild animals: deer, frog, horse. Furthermore, for the case with $G=2$, we have the groups/superclasses animals, and no-animals, as in bai2021clustering.
Next, we present additional results for CIFAR-100 (and MobileNet) in Table 9, where we consider 20 superclasses (of labels) as in ramasesh2020anatomy. We remind the reader that, in Table 4 in the main text, we saw that there is a trade-off between memory reduction and performance when using fewer heads. Even though there is a slight decrease in performance as we use fewer heads, SFL+Hydra still outperforms SFL. Hence, with Hydra, we can reduce the memory demands on the server while still having a better performance than baseline SFL. On the other hand, in Table 9, for CIFAR-100, we see that the two configurations with $G=20$ and $G=L=100$ have similar performance. This can be justified by an implicit observation derived from Table 1 in Section 3.2: datasets with a larger number of labels are less sensitive to data heterogeneity and CF in SFL than datasets with a smaller number of labels. In particular, in Table 1, we see that the drop in accuracy as the data heterogeneity increases (compared to IID) is larger in CIFAR-10 than in CIFAR-100. It is worth noting that similar observations have been made by related work on FL, e.g., see lee2022preservation. Therefore, the fact that the variations in PG and global accuracy are smaller in Table 9 (for CIFAR-100) than the ones in Table 4 (for CIFAR-10) is related to the observation above.
@@
Dataset Method Number of Performance Gap $\downarrow$ Global Accuracy $\uparrow$

### v2

@@ E.1 Dominant Label Ratio
E.2 Dirichlet Partitioning Method
Dirichlet partitioning is probably the most commonly used method to federate data in distributed computing research (Jimenez G. et al., 2024). It leverages the Dirichlet Distribution (Minka, 2000) to allocate samples among clients, effectively simulating non-IID scenarios. The concentration parameter, $\alpha$, determines the level of data heterogeneity: lower values result in more imbalanced distributions, while higher values produce distributions closer to IID conditions. This approach enables the creation of diverse non-IID datasets, facilitating the evaluation of FL algorithms under varying levels of data imbalance.
@@ E.3 Custom Sharding Method
The sharding technique (Zhang et al., 2021) generates non-IID datasets by first sorting data based on class labels and dividing it into smaller, equally sized subsets known as shards. These shards are then distributed to clients, with each client receiving a fixed number. As each shard contains data from only a few classes, clients end up with datasets representing only a limited portion of the overall class distribution. This approach creates a highly skewed, non-IID distribution, where clients predominantly have samples from a small number of classes, without any overlap in class representation across clients from different shards.
@@ F.2 Cyclic-and-reverse Order
Appendix G Extended Systematic Analysis of Catastrophic Forgetting in SFL
@@
G.1 Performance under Other Data Partitions
Table 5: Global accuracy (reported median of the last five rounds across all runs) and PG of SFL for other data partitions, supplementary to Table 1. The results presented here are for cyclic order with $\phi=10$. The parentheses present the global accuracy and PG, respectively, for IID.
Global Accuracy $\uparrow$ Performance Gap $\downarrow$
@@ G.1 Performance under Other Data Partitions
CIFAR-10 $(\mathbf{K=10})$
MobileNet (80, 10) $\mathbf{50}$ $\mathbf{\pm 0.4}$ 50 $\mathbf{\pm 0.5}$ $40$ $\pm 0.6$ 36 $\mathbf{\pm 1}$ $43.7$ $\pm 0.5$ $45.5$ $\pm 1$
ReNet101 (68, 15) 42 $\mathbf{\pm 1.2}$ $33$ $\pm 0.6$ $27$ $\pm 1$ 35 $\mathbf{\pm 3}$ $53$ $\pm 2$ $55$ $\pm 3$
TinyImageNet $(\mathbf{K=200})$
MobileNet (50, 40) $27$ $\pm 0.3$ 45 $\mathbf{\pm 0.3}$ $31$ $\pm 1$ $62$ $\pm 1.6$ 46 $\mathbf{\pm 1}$ $63$ $\pm 0.7$
@@ G.4 Studying the Per-Position Accuracy
Per-position Accuracy. The position denotes the point in the processing order at which client $c_{l}$, whose dominant label is $l$, is served by the server. In practice, the position-$k$ accuracy (for $k\in\{1\ldotp\ldotp L\}$,) corresponds to the accuracy of the labels that are dominant among the clients processed at position $k$ within each processing cycle of length $L$. To ensure that the results are independent of the type of label, the order of the labels is randomly allocated in every experiment, and then the median value from all repeated experiments is computed to give the per-position accuracy. This definition is tailored to the cyclic order and DL partitioning, but can be extended to other types, too.
In the following, we present supplemental results to the ones presented in Section 3. Note that we incorporate three different models, and for each of them, we use three different cut-layer levels (i.e., shallow-cut, mid-cut, deep-cut). In detail, the definition of these cuts, respectively, for each level, is: (i) MobileNet has $26$ layers in total and uses cuts $4,15$, and $23$, (ii) ResNet101 has $37$ layers in total and uses cuts $2,10$, and $33$, (iii) ResNet18 has $12$ layers in total and uses cuts $2,6$, and $10$. Note that for the ResNet models we consider that each residual block is a single layer, like in (Tirana et al., 2024). Moreover, recall that the cut-layer selection follows the bottleneck policy as discussed in Appendix D.2.
@@
G.5 Impact of the Cut Layer and Comparison with SplitFedv1/FL
@@
Figure 16 presents an extended version of the scatter plot shown in Figure 3, which depicts the global accuracy and the PG for different cut layers (in SFL), and for SplitFedv1 (Thapa et al., 2022), a variant of SL. It has been shown that its performance is the same as that of FL (Thapa et al., 2022), and hence, we denote it by FL. Here we utilize the same cut-layers as in Appendix G.4. We see that in terms of accuracy, SFL performs better than FL under IID data. Still, as data heterogeneity increases, FL performs better. While these results concern cyclic order and DL partition, results on random order and Dirichlet partition are provided in (Han et al., 2024). Similar are the results for CIFAR-100, even though the performance is not affected as much when the data heterogeneity increases. This is a side-effect of our earlier observation made in Appendix G.4 (i.e., lower $L$, present higher intraCF).
@@ Appendix H Extended Numerical Evaluation
Global Accuracy $\uparrow$ Performance gap $\downarrow$
@@
SFL $7.4_{\pm 0.4} 77$$7.7_{\pm 0.4}$ $8.2_{\pm 0.5} 77$$7.7_{\pm 0.6}$ $24_{\pm 0.3}$ $24_{\pm 0.4}$ $21.8_{\pm 0.2}$ $23.8_{\pm 0.7}$
SFL+Hydra $\mathbf{81.3_{\pm 0.1}}$ $\mathbf{81_{\pm 0.1}}$ $\mathbf{81_{\pm 0.1}}$ $\mathbf{82_{\pm 0.3}}$ $\mathbf{8.3_{\pm 0.6}}$ $\mathbf{8.3_{\pm 0.5}}$ $\mathbf{10_{\pm 0.4}}$ $\mathbf{9.3_{\pm 0.6}}$
@@
SFL $56_{\pm 5}$ $7_{\pm 3} 66$$6.2_{\pm 1.6}$ $59.7_{\pm 2}$ $50_{\pm 3}$ $46_{\pm 3}$ $42_{\pm 1}$ $54.9_{\pm 1}$
SFL+Hydra $\mathbf{75.2_{\pm 1.4}}$ $\mathbf{74.6_{\pm 0.8}}$ $\mathbf{73.4_{\pm 1.5}}$ $\mathbf{66_{\pm 0.9}}$ $\mathbf{13_{\pm 0.9}}$ $\mathbf{13_{\pm 1.5}}$ $\mathbf{15_{\pm 0.6}}$ $\mathbf{12_{\pm 0.8}}$
@@
Global Accuracy $\uparrow$ Performance Gap $\downarrow$
@@
SFL+Hydra $\mathbf{45_{\pm 1}}$ $\mathbf{45.7_{\pm 1}}$ $\mathbf{39_{\pm 0.4}}$ $\mathbf{22_{\pm 2}}$ $\mathbf{25_{\pm 2}}$ $\mathbf{30_{\pm 2}}$
SFL $42_{\pm 1.2}$ $33_{\pm 0.6}$ $27_{\pm 1}$ $35_{\pm 3}$ $3_{\pm 2} 55$$5_{\pm 3}$
@@ H.1 SFL and Hydra
H.2 Results with Other Types of Ordering
@@ H.3 Results with Other Data Partitioning Methods
Global Accuracy $\uparrow$ Performance Gap $\downarrow$
@@
SFL+Hydra $\mathbf{66.5}$$\mathbf{\pm 0.1}$ $\mathbf{74.6{\pm 0.1}}$ $\mathbf{45_{\pm 0.1}}$ $\mathbf{13.4_{\pm 1}}$ $\mathbf{12_{\pm 1}}$ $\mathbf{26.3_{\pm 0.66}}$
SplitFedV1/FL $55_{\pm 0.2}$ $63.4_{\pm 0.1}$ $32_{\pm 0.19}$ $15.7_{\pm 1.5}$ $15{\pm 0.4}$ $34.46_{\pm 0.3}$
SplitFedV3 $43_{\pm 0.3}$ $51.4_{\pm 0.1}$ $5.9_{\pm 0.1} 22$$6_{\pm 0.8} 22$$2_{\pm 1}$ $47.3_{\pm 0.6}$
MultiHead $47_{\pm 0.3}$ $55.7_{\pm 0.46}$ $25.4_{\pm 0.18}$ $36.6_{\pm 0.9}$ $30.5_{\pm 1}$ $59.9_{\pm 0.3}$
@@
SFL + Hydra $\mathbf{49.7_{\pm 1.4}}$ $\mathbf{51.9_{\pm 0.9}}$ $\mathbf{33_{\pm 0.9}}$ $\mathbf{16.6_{\pm 3.4}}$ $\mathbf{17_{\pm 2.4}}$ $\mathbf{33_{\pm 1}}$
SplitFedV1/FL $29.7_{\pm 1.8}$ $39_{\pm 1.1}$ $3_{\pm 1.6} 22$$2.8_{\pm 2.4}$ $24.1_{\pm 4}$ $43_{\pm 1.6}$
SplitFedV3 $19.1_{\pm 1.6}$ $35_{\pm 2}$ $40_{\pm 3}$ $27.1_{\pm 1}$
MultiHead $29_{\pm 1}$ $40_{\pm 0.9}$ $33_{\pm 2}$ $28_{\pm 4}$
H.4 Comparison of Hydra with Baselines and State-of-the-Art Methods
@@ H.5 Weight Regularization
H.6 Ablation Study on the Number of Heads ($G$)
In this section, we first provide details about the definition of the superclasses in CIFAR-10 for the results presented in Table 4. In particular, similar in spirit to (Bai et al., 2021), we group labels according to their semantic relationships. In detail, for $G=5$, we have defined the groups/superclasses: land vehicles: automobile, truck, things found in the sky: airplane, bird, water vehicles: ship, domestic animals: cat, dog, wild animals: deer, frog, horse. Furthermore, for the case where $G=2$, we have defined the groups/superclasses animals, and no-animals, as in (Bai et al., 2021).
Next, we present additional results for CIFAR-100 (trained with MobileNet) in Table 9, where we consider 20 superclasses (of labels) as in (Ramasesh et al., 2021a). We remind the reader that, in Table 4 in the main text, we saw that there is a trade-off between memory reduction and performance when using fewer heads. In particular, even though there is a slight decrease in performance as we use fewer heads, SFL+Hydra still outperforms SFL. Hence, with Hydra, we can reduce the memory demands on the server while still having a better performance than the baseline SFL. On the other hand, in Table 9, for CIFAR-100, we see that the two configurations with $G=20$ and $G=L=100$ have similar performance. This can be justified by an implicit observation derived from Table 1 in Section 3.2: datasets with a larger number of labels are less sensitive to data heterogeneity and CF in SFL than datasets with a smaller number of labels. In particular, in Table 1, we see that the drop in accuracy as the data heterogeneity increases (compared to IID) is larger in CIFAR-10 than in CIFAR-100. It is worth noting that similar observations have been made by related work on FL, e.g., see (Lee et al., 2022). Therefore, the fact that the variations in PG and global accuracy are smaller in Table 9 (for CIFAR-100) than the ones in Table 4 (for CIFAR-10) may be related to the observation above.
@@ H.6 Ablation Study on the Number of Heads ($G$)
Dataset Method Number of Global Accuracy $\uparrow$ Performance Gap $\downarrow$